OpenStudy (anonymous):

can some one help me. on this math problem thank you!! Given f(x) = the quantity of 4x plus 1, divided by 3 solve for f-1(3).

OpenStudy (bibby):

how would you get the inverse of \[y = \frac{ 4x+1 }{ 3 }\]

OpenStudy (bibby):

here's a hint on how to get started on finding the inverse\[x = \frac{ 4y+1 }{ 3 }\]

OpenStudy (anonymous):

well probably change were is 3 you put it on were is 4 will be like x=(3y+1) /4

OpenStudy (bibby):

You just solve for y as you would for any equation only this y is for f-1(x) and not f(x) if you multiply both sides by 3 and are still solving for y:\[3x = 4y+1\]

OpenStudy (anonymous):

so I subtract 4 to both sides

OpenStudy (bibby):

how can you subtract 4? "4y" is the whole term. you can subtract 4y but we don't want to do that. We want to get y alone. however if we subtract 1\[4y=3x-1\]

OpenStudy (anonymous):

so then I should divide by 4 to live the y alone.

OpenStudy (bibby):

yeah and that is the formula for f-1(x)

OpenStudy (anonymous):

ok so which of this will be the answer A) 2 B)4 C)6 D)8

OpenStudy (anonymous):

so the D will be the answers

OpenStudy (bibby):

I don't know. Why don't you plug in 3 into \[f^-1(x) = \frac{ 3x-1 }{ 4 }\]

OpenStudy (anonymous):

so it will be A can you check that out please!

OpenStudy (bibby):

yeah that's right

OpenStudy (anonymous):

thank you! i just want to be sure